OpenStudy (anonymous):

If the range of the function f(x) = 4x – 3 is {11.4, 15, 17, 29}, what is its domain?

OpenStudy (anonymous):

A. {1.5, 2.8, 5.5, 7} B. {2.3, 4, 8, 11} C. {7, 11, 13.2, 14} D. {3.4, 5, 8, 9.3} E. {3.6, 4.5, 5, 8}

OpenStudy (anonymous):

???

OpenStudy (jdoe0001):

hmmm I see hehe

OpenStudy (anonymous):

Lol do you know the answer please?

OpenStudy (ivettef365):

you need to find f(x) when x = 11, 4, 15 .......

OpenStudy (jdoe0001):

what is really saying is that in f(x) = 4x – 3 ==> y = = 4x – 3 if "y" equals all these values below, one at a time {11.4, 15, 17, 29} what will be the value of "x" so say let's pick the 1st 2 values "y = 11" 11 = 4x -3 # solving for "x" 14 = 4x => 14/4 = x now if "y = 4" 4 = 4x -3 # solving for "x" 7 = 4x => 7/4 = x

OpenStudy (anonymous):

Can it be plugged into a graphing calculator?

OpenStudy (jdoe0001):

sure

OpenStudy (jdoe0001):

but I think you're expected to solve for "x" anyway